OpenStudy (anonymous):

How do I graph the equations y=3x+8 and y=4x+5?

OpenStudy (amistre64):

graph? or sketch?

OpenStudy (anonymous):

Graph

OpenStudy (amistre64):

if sketch (a rough resemblence with minimal markers) then you would find the intercepts

OpenStudy (amistre64):

determine the value y when, and then the value of x when y=0 this give syou 2 points to plot, then draw a line thru the points

OpenStudy (amistre64):

or, plot the y intercept that is the ending number, and follow the slope defined by the coefficient of the x part

OpenStudy (amistre64):

or, determine 2 points of reference by trial and use them to draw a line thru ... there are many ways to approach this

OpenStudy (anonymous):

hmmm

OpenStudy (anonymous):

It just wants me to put both lines on a graph

OpenStudy (anonymous):

Im not sure what that means, its just hat the instructions say

jimthompson5910 (jim_thompson5910):

Graphing y=3x+8 This is a linear equation which graphs a straight line. So we just need 2 points If x = 0, then y=3x+8 y=3*0+8 y = 8 So if x = 0, then y = 8 giving us the ordered pair (0,8) which is one point on the line If x = 1, then y=3x+8 y=3*1+8 y = 11 So we know the point (1,11) is also another point on the line Plot the two points (0,8) and (1,11) and draw a straight line through them. Make sure this line extends infinitely in both directions.

jimthompson5910 (jim_thompson5910):

you will follow the same steps when it comes to graphing y=4x+5

OpenStudy (anonymous):

thank you!

OpenStudy (rational):

step1 : plot both the points (0, 8) and (1, 11) 

OpenStudy (rational):

step2 : draw a line passing through those two points